# Beyond Cascaded Architectures: An End-to-end Generative Framework for Industrial Advertising

### v2

@@ Abstract.
To bridge this gap, we introduce End-to-End Generative Advertising (EGA), the first unified framework that holistically models user interests, point-of-interest (POI) and creative generation, ad allocation, and payment optimization within a single generative model. Our approach employs hierarchical tokenization and multi-token prediction to jointly generate POI recommendations and ad creatives, while a permutation-aware reward model and token-level bidding strategy ensure alignment with both user experiences and advertiser objectives. Additionally, we decouple allocation from payment using a differentiable ex-post regret minimization mechanism, guaranteeing approximate incentive compatibility at the POI level. Through extensive offline evaluations and large-scale online experiments on real-world advertising platforms, we demonstrate that EGA significantly outperforms traditional cascaded systems in both performance and practicality. Our results highlight its potential as a pioneering fully generative advertising solution, paving the way for next-generation industrial ad systems.
@@ 1. Introduction
To address above challenges, we propose End-to-end Generative Advertising (EGA), a novel generative framework that unifies all decision-making stages into a single model. EGA bridge the gap between generative modeling and the practical requirements of industrial advertising, which directly outputs the final ad sequence, as well as corresponding creatives, positions, and payment end-to-end. Firstly, inspired by generative recommendation techniques, we leverage Residual Quantized Variational AutoEncoder (RQ-VAE) to encode user behavior and item features into hierarchical semantic tokens, and employ an encoder-decoder architecture with multi-token prediction to jointly generate candidate ad sequences and creative content. Secondly, we propose a token-level bidding and generative allocation mechanism, enhanced by permutation-aware reward modeling. This strategy decouples allocation from payment, allowing the model to effectively reflect business objectives during generation while approximately preserving incentive compatibility (IC) with a differentiable payment network. Last but not least, we introduce a multi-phase training paradigm. A pre-training phase learns user interests from actual exposure sequences that contain ad and organic items. Then auction-based post-training is applied to fine-tune the model with ad-specific constraints, dynamically optimizing ad allocation based on auction signals and platform objectives.
@@
  • We introduce EGA, which surpasses traditional multi-stage cascading architectures by an unified single generative model. To the best of our knowledge, this is one of the first end-to-end generative advertising framework in industry.
@@
  • Extensive offline experiments and online A/B tests in large-scale industrial datasets demonstrate the effectiveness and efficiency of our approach.
@@ 4. Methodology
We introduce EGA, an end-to-end generative advertising framework. This section describes how we learn user interests via vector tokenization and encoder-decoder design, leverage permutation-aware reward modeling for business objectives, incorporate auction-based preference alignment, and integrate all components through a unified multi-phase optimization strategy.
@@ 4.1.3. Probabilistic Decomposition
The modeling of the target item’s probability distribution is decomposed into two stages, leveraging the complementary advantages of POI-level and creative-level representations. Rather than directly predicting the next item $\bm{a}_{t+1}$ from the historical interaction sequence $\mathcal{S}^{u}_{1:t}$, EGA first predicts the POI $\bm{a}_{t+1}^{poi}$, then determines the creative image $\bm{a}_{t+1}^{img}$.
@@ 4.4. Optimization and Training
Figure 4. Overview of the proposed optimization and training pipeline. The training process consists of two steps: 1) Interest-based Pre-training, where behavior sequences are used to pre-train EGA that predicts the next POI and creative tokens; and 2) Auction-based Post-training, which includes: 2.1) a permutation-aware reward model trained with list-wise feedback to predict reward signals, such as pCTR and pCVR; 2.2) generative allocation, where different candidate sequences are scored by the reward model to optimize policy gradients; and 2.3) a payment network optimized with ex-post regret to ensure IC approximately.
@@ 5. Experiments
In this section, we evaluate our proposed model on industrial dataset and aim to answer the following research questions:
  • RQ1: How does our EGA model perform, compared to the state-of-the-art advertising models?
  • RQ2: What is the impact of designs (e.g. MTP module, token-level bidding, payment network, and multi-phase training) on the performance of EGA?
@@
  • RQ4: How does EGA perform in the real-world industrial advertising scenarios?
@@ 5.1.1. Dataset
The industrial dataset used in our experiments consists of real interaction logs collected from a large-scale location-based services (LBS) platform, spanning the period from September 2024 to April 2025. The dataset contains 200 million requests from over 2 million users and nearly 10 million unique ads. We use the first 200 days for pre-training and randomly sample 10% data for preference alignment. The last 14 days are used for testing.
@@ 5.1.3. Baselines
We evaluate EGA against the following two widely adopted industrial architectures.
@@ 5.1.4. Implementation Details
We train EGA using the Adam optimizer with an initial learning rate of 0.0024. The batch size is set to 128. Model training and optimization are performed on NVIDIA A100 GPUs with 80G memory. For hyperparameters, we tried different hyperparameters using grid search. Due to space constraints, we report only the most optimal hyperparameter settings in this paper. For interest-based pre-training, the block number $L$ of encoder and decoder is 3, the number of codebook layers $C=3$, and codebook size $W$ of each layer is 1024. We consider $K=10$ target session items and use $B=256$ historical behaviors as context. For auction-based preference alignment, the hyperparameters in token-level bidding $\alpha=1.2$ and $\beta=2$. We generate $N_{S}=64$ different sequences for each request by beam search. For reward model and payment network, the hidden layers of the MLP are 128, 32, and 10.
@@ 5.2. Offline Performance (RQ1)
Table 2. The experimental results of out model EGA and competitors on industrial dataset. The bold value marks the best one in each column. Each result is presented in the form of mean (lift percentage). Lift percentage means the improvement of EGA over the best baselines.
@@
Our proposed EGA overcomes these shortcomings by integrating end-to-end generation, permutation-aware reward modeling, token-level bidding, and a dedicated payment network. As a result, EGA achieves the best overall performance: it significantly improves revenue (RPM), enhances both POI and creative CTRs, and delivers superior economic robustness with the lowest IC regret among all baselines. This demonstrates that addressing both early-stage candidate loss and business constraints is crucial for practical deployment in industrial advertising systems.
@@ 5.3. Ablation Study (RQ2)
Table 3. Ablation Study of EGA.
@@
We conduct experiments to validate the effectiveness of different components. Correspondingly, we design a series of ablation studies, which considers four variants to simplify EGA in different ways:
@@
The performance results are shown in Table 3, from which we observe the full EGA model consistently outperforms all ablation variants across all major metrics, validating the effectiveness of our design. Specifically,
@@
In summary, each component of EGA, including multi-token prediction, multi-phase training, max aggregation in bidding, and a dedicated payment network, plays an important and complementary role in improving revenue and user experience, and ensuring economic robustness.
@@ 5.4. Hyperparameters (RQ3)
Figure 5. Effect of hyperparameters of EGA.
We conduct a comprehensive study to evaluate the sensitivity of EGA with respect to key hyperparameters. The results are summarized in Figure 5, where RPM and CTR-poi are used as the evaluation metrics. Figure 5 examines the impact of token-level bidding parameters $\alpha$ and $\beta$ in Equation (21). $\alpha$ measures the importance of bids during the allocation process. Increasing $\alpha$ significantly boosts RPM at first, as higher-bid ads become more likely to win in the generative allocation. However, this trend gradually plateaus as further increasing $\alpha$ provides diminishing marginal returns. In contrast, CTR-poi shows a consistent decline as $\alpha$ rises, reflecting a shift in exposure from more organic contents to higher-bid ads. Besides, $\beta$ acts as a global balancing factor between ads and organic content. Increasing $\beta$ reduces the probability of ads being allocated. As a result, RPM decreases steadily with increasing $\beta$, and eventually flattens out as ads rarely win positions in the list. Conversely, CTR-poi increases with $\beta$, as the display list becomes dominated by organic content. In practice, the optimal values of $\alpha$ and $\beta$ are selected to achieve a trade-off between platform revenue and user experience.
5.5. Online A/B Test (RQ4)
To validate the real-world effectiveness of our proposed EGA framework, we conducted online A/B tests on an industrial advertising platform, comparing EGA with MCA baseline. Table 4 summarizes the results of online A/B testing conducted from April 2 to April 8, 2025, utilizing 2% of total production traffic. EGA achieves a 15.2% increase in RPM, a 6.4% improvement in CTR, and a 3.1% lift in return on investment (ROI), compared to the MCA baseline. These consistent improvements across revenue, user experience, and advertiser utility highlight the practical value of EGA in a complex and production-scale environment. Besides, benefiting from algorithm engineering and system optimizations, the online response time (RT) of EGA increases by only 7 ms per request (2.5% relatively). This marginal overhead demonstrates that EGA maintains high computational efficiency.
Table 4. The experimental results from Online A/B tests.
Relative change in metrics RPM CTR ROI RT
EGA over baseline-MCA +15.2% +6.4% +3.1% +2.5%
@@ 6. Conclusion
In this work, we presented End-to-end Generative Advertising (EGA), a novel framework that unifies ranking, creative selection, ad allocation, and payment computation into a single generative model for industrial advertising systems. By leveraging hierarchical semantic tokenization, permutation-aware reward modeling, and token-level bidding and allocation, EGA bridges the gap between user interests modeling and business-critical auction constraints such as IC and IR. The proposed multi-phase training paradigm, including interest-based pre-training and auction-based post-training, ensures that EGA captures both user interests and advertiser utility under complex real-world conditions. Extensive offline experiments and large-scale online A/B tests on an industrial dataset demonstrate that EGA achieves substantial improvements over both multi-stage cascading architecture and recent generative recommendation baselines in platform revenue, user experience, and advertiser return on investment, all with minimal latency overhead. Our ablation studies validate the effectiveness of each design component and highlight the flexibility of the proposed framework.
Looking forward, we believe EGA opens new directions for the integration of generative modeling and economic mechanism design in online advertising. Future work will further explore scaling laws and enhance business interpretability.

### v3

# EGA-V2: An End-to-end Generative Framework for Industrial Advertising

@@ Abstract.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@
To bridge this gap, we introduce End-to-End Generative Advertising (EGA-V2), the first unified framework that holistically models user interests, point-of-interest (POI) and creative generation, ad allocation, and payment optimization within a single generative model. Our approach employs hierarchical tokenization and multi-token prediction to jointly generate POI recommendations and ad creatives, while a permutation-aware reward model and token-level bidding strategy ensure alignment with both user experiences and advertiser objectives. Additionally, we decouple allocation from payment using a differentiable ex-post regret minimization mechanism, guaranteeing approximate incentive compatibility at the POI level. Through extensive offline evaluations we demonstrate that EGA-V2 significantly outperforms traditional cascaded systems in both performance and practicality. Our results highlight its potential as a pioneering fully generative advertising solution, paving the way for next-generation industrial ad systems.
@@ 1. Introduction
To address above challenges, we propose End-to-end Generative Advertising (EGA-V2), a novel generative framework that unifies all decision-making stages into a single model. EGA-V2 bridge the gap between generative modeling and the practical requirements of industrial advertising, which directly outputs the final ad sequence, as well as corresponding creatives, positions, and payment end-to-end. Firstly, inspired by generative recommendation techniques, we leverage Residual Quantized Variational AutoEncoder (RQ-VAE) to encode user behavior and item features into hierarchical semantic tokens, and employ an encoder-decoder architecture with multi-token prediction to jointly generate candidate ad sequences and creative content. Secondly, we propose a token-level bidding and generative allocation mechanism, enhanced by permutation-aware reward modeling. This strategy decouples allocation from payment, allowing the model to effectively reflect business objectives during generation while approximately preserving incentive compatibility (IC) with a differentiable payment network. Last but not least, we introduce a multi-phase training paradigm. A pre-training phase learns user interests from actual exposure sequences that contain ad and organic items. Then auction-based post-training is applied to fine-tune the model with ad-specific constraints, dynamically optimizing ad allocation based on auction signals and platform objectives.
@@
  • We introduce EGA-V2, which surpasses traditional multi-stage cascading architectures by an unified single generative model. To the best of our knowledge, this is one of the first end-to-end generative advertising framework in industry.
@@
  • Extensive offline experiments in large-scale industrial datasets demonstrate the effectiveness and efficiency of our approach.
@@ 4. Methodology
We introduce EGA-V2, an end-to-end generative advertising framework. This section describes how we learn user interests via vector tokenization and encoder-decoder design, leverage permutation-aware reward modeling for business objectives, incorporate auction-based preference alignment, and integrate all components through a unified multi-phase optimization strategy.
@@ 4.1.3. Probabilistic Decomposition
The modeling of the target item’s probability distribution is decomposed into two stages, leveraging the complementary advantages of POI-level and creative-level representations. Rather than directly predicting the next item $\bm{a}_{t+1}$ from the historical interaction sequence $\mathcal{S}^{u}_{1:t}$, EGA-V2 first predicts the POI $\bm{a}_{t+1}^{poi}$, then determines the creative image $\bm{a}_{t+1}^{img}$.
@@ 4.4. Optimization and Training
Figure 4. Overview of the proposed optimization and training pipeline. The training process consists of two steps: 1) Interest-based Pre-training, where behavior sequences are used to pre-train EGA-V2 that predicts the next POI and creative tokens; and 2) Auction-based Post-training, which includes: 2.1) a permutation-aware reward model trained with list-wise feedback to predict reward signals, such as pCTR and pCVR; 2.2) generative allocation, where different candidate sequences are scored by the reward model to optimize policy gradients; and 2.3) a payment network optimized with ex-post regret to ensure IC approximately.
@@ 5. Experiments
In this section, we evaluate our proposed model on industrial dataset and aim to answer the following research questions [^2] Due to computational complexity constraints, our current evaluation is limited to offline experiments, with online A/B testing reserved for future work.:
  • RQ1: How does our EGA-V2 model perform, compared to the state-of-the-art advertising models?
  • RQ2: What is the impact of designs (e.g. MTP module, token-level bidding, payment network, and multi-phase training) on the performance of EGA-V2?
@@ 5.1.1. Dataset
The industrial dataset used in our experiments consists of real interaction logs collected from a large-scale location-based services (LBS) platform Meituan, spanning the period from September 2024 to April 2025. The dataset contains 200 million requests from over 2 million users and nearly 10 million unique ads of low-traffic ad slot in Meituan. We use the first 200 days for pre-training and randomly sample 10% data for preference alignment. The last 14 days are used for testing.
@@ 5.1.3. Baselines
We evaluate EGA-V2 against the following two widely adopted industrial architectures.
@@ 5.1.4. Implementation Details
We train EGA-V2 using the Adam optimizer with an initial learning rate of 0.0024. The batch size is set to 128. Model training and optimization are performed on NVIDIA A100 GPUs with 80G memory. For hyperparameters, we tried different hyperparameters using grid search. Due to space constraints, we report only the most optimal hyperparameter settings in this paper. For interest-based pre-training, the block number $L$ of encoder and decoder is 3, the number of codebook layers $C=3$, and codebook size $W$ of each layer is 1024. We consider $K=10$ target session items and use $B=256$ historical behaviors as context. For auction-based preference alignment, the hyperparameters in token-level bidding $\alpha=1.2$ and $\beta=2$. We generate $N_{S}=64$ different sequences for each request by beam search. For reward model and payment network, the hidden layers of the MLP are 128, 32, and 10.
@@ 5.2. Offline Performance (RQ1)
Table 2. The experimental results of out model EGA-V2 and competitors on industrial dataset. The bold value marks the best one in each column. Each result is presented in the form of mean (lift percentage). Lift percentage means the improvement of EGA-V2 over the best baselines.
@@
Our proposed EGA-V2 overcomes these shortcomings by integrating end-to-end generation, permutation-aware reward modeling, token-level bidding, and a dedicated payment network. As a result, EGA-V2 achieves the best overall performance: it significantly improves revenue (RPM), enhances both POI and creative CTRs, and delivers superior economic robustness with the lowest IC regret among all baselines. This demonstrates that addressing both early-stage candidate loss and business constraints is crucial for practical deployment in industrial advertising systems.
@@ 5.3. Ablation Study (RQ2)
Table 3. Ablation Study of EGA-V2.
@@
We conduct experiments to validate the effectiveness of different components. Correspondingly, we design a series of ablation studies, which considers four variants to simplify EGA-V2 in different ways:
@@
The performance results are shown in Table 3, from which we observe the full EGA-V2 model consistently outperforms all ablation variants across all major metrics, validating the effectiveness of our design. Specifically,
@@
In summary, each component of EGA-V2, including multi-token prediction, multi-phase training, max aggregation in bidding, and a dedicated payment network, plays an important and complementary role in improving revenue and user experience, and ensuring economic robustness.
@@ 5.4. Hyperparameters (RQ3)
Figure 5. Effect of hyperparameters of EGA-V2.
We conduct a comprehensive study to evaluate the sensitivity of EGA-V2 with respect to key hyperparameters. The results are summarized in Figure 5, where RPM and CTR-poi are used as the evaluation metrics. Figure 5 examines the impact of token-level bidding parameters $\alpha$ and $\beta$ in Equation (21). $\alpha$ measures the importance of bids during the allocation process. Increasing $\alpha$ significantly boosts RPM at first, as higher-bid ads become more likely to win in the generative allocation. However, this trend gradually plateaus as further increasing $\alpha$ provides diminishing marginal returns. In contrast, CTR-poi shows a consistent decline as $\alpha$ rises, reflecting a shift in exposure from more organic contents to higher-bid ads. Besides, $\beta$ acts as a global balancing factor between ads and organic content. Increasing $\beta$ reduces the probability of ads being allocated. As a result, RPM decreases steadily with increasing $\beta$, and eventually flattens out as ads rarely win positions in the list. Conversely, CTR-poi increases with $\beta$, as the display list becomes dominated by organic content. In practice, the optimal values of $\alpha$ and $\beta$ are selected to achieve a trade-off between platform revenue and user experience.
@@ 6. Conclusion
In this work, we presented End-to-end Generative Advertising (EGA-V2), a novel framework that unifies ranking, creative selection, ad allocation, and payment computation into a single generative model for industrial advertising systems. By leveraging hierarchical semantic tokenization, permutation-aware reward modeling, and token-level bidding and allocation, EGA-V2 bridges the gap between user interests modeling and business-critical auction constraints such as IC and IR. The proposed multi-phase training paradigm, including interest-based pre-training and auction-based post-training, ensures that EGA-V2 captures both user interests and advertiser utility under complex real-world conditions. Extensive offline experiments on an industrial dataset demonstrate that EGA-V2 achieves substantial improvements over both multi-stage cascading architecture and recent generative recommendation baselines in platform revenue, user experience, and advertiser return on investment. Our ablation studies validate the effectiveness of each design component and highlight the flexibility of the proposed framework.
Looking forward, we believe EGA-V2 opens new directions for the integration of generative modeling and economic mechanism design in online advertising. Future work will further explore scaling laws and enhance business interpretability.